OpenStudy (anonymous):

what would be a graph of y equals cot x. also please explain why it would be that graph.

OpenStudy (australopithecus):

y = cot(x) cot = 1/tan = cos/sin where sin(x) =/= 0 because you cannot divide by zero thus we will have vertical asymptotes anywhere in the domain where sin(x) = 0 the graph would appear furthermore since